OpenStudy (theghostlycreeper):

What is the value of x? 27(x+4)=−627(x+4)=−6 −35−35 −25−25 25 35

OpenStudy (theghostlycreeper):

@KJSaif

OpenStudy (theghostlycreeper):

ya alones

alones (alones):

Huh?

OpenStudy (theghostlycreeper):

just saying ya

alones (alones):

Alright so distribute at first \(\ (27)(x)+(27)(4)=−6\) and this would look like this \(\ 27x+108=−6\) as simplifyed

alones (alones):

Alright so than subtract\(108\) from each side \[\ 27x+108−108=−6−108\] \[\ 27x=−114\]

alones (alones):

At last divide both sides by \(27.\) \[\frac{ 27x }{ 27 }=\frac{ −114 }{ 27 }\]

alones (alones):

@theghostlycreeper divide \[\frac{ −114 }{ 27 }\] and tell me what you got.

OpenStudy (theghostlycreeper):

-25

OpenStudy (theghostlycreeper):

was that right